# Cleaning Up the Streets: Understanding Motivations, Mental Models, and Concerns of Users Flagging Social Media Posts

### v2

@@ Abstract.
Social media platforms offer flagging, a technical feature that empowers users to report inappropriate posts or bad actors to reduce online harm. The deceptively simple flagging interfaces on nearly all major social media platforms disguise complex underlying interactions among users, algorithms, and moderators. Through semi-structured interviews with 25 social media users experienced in flagging inappropriate content, we examine end-users’ understanding of flagging procedures, explore the factors that motivate them to flag, and surface their cognitive and privacy concerns. Our findings uncover a lack of procedural transparency in flagging mechanisms that create gaps in users’ mental models. Regardless, users strongly believe that it is crucial for platforms to provide flagging options. We highlight how flags raise questions about distributing labor and responsibility between platforms and users for addressing online harm. We recommend innovations in the flagging design space that assist user comprehension, ensure privacy, and reduce cognitive burdens.
@@ 1. Introduction
End-users have several options to take action when encountering inappropriate content on social media platforms, from regular interactions built into a platform (e.g., unfollowing an account) to content moderation-specific interactions. Of these options, it is only through flagging that end-users can directly notify platform administrators. Thus, the flag serves as a powerful and empowering tool for user-driven moderation.
Flagging a post involves three temporally distinct stages from the flagger’s perspective: (1) before, (2) during, and (3) after flagging. Each step grows more complex upon closer inspection. How does a user decide if the content is inappropriate? What do they do when platform guidelines are unclear? How do users interact with flag submission interfaces? What makes a flag submission effective, ensuring users feel their efforts weren’t wasted? In asking these questions, we are led to Crawford and Gillespie’s description of flagging as a “a complex interplay between users and platforms, humans and algorithms, and the social norms and regulatory structures of social media”  (Crawford and Gillespie, 2016). It is with this in mind that we examine the perspectives of users through these three steps of the flagging processes, ultimately aiming to determine how we may better understand flagging mechanisms to improve them for users.
@@
Though at first glance, flag implementations may signal platforms’ commitment to enacting democratic governance and fostering user communities, in practice, flags lack transparency and are often designed without adequately considering users’ needs. Most major platforms that host user-generated content, e.g., Instagram (Fig. 1), thank users for flagging content, tell them they will be notified when a decision is made, and provide them with additional safety options such as blocking, muting, and hiding the flagged content. However, only Facebook offers users a centralized interface where they may revisit past flags to check their review status (Fig. 2).
(a) A thank you message to users for flagging a post on Instagram.
(b) Additional options to block and hide content after flagging on Instagram.
Figure 1. Instagram’s flagging interface pages.
(a) A list of submitted reports, as listed in Facebook’s support inbox.
(b) A thank you message to users for flagging a post on Facebook.
Figure 2. Facebook’s flagging interface pages.
@@
In this study, we analyze user experiences with flagging mechanisms on major social media platforms and ask three key research questions. First, given the limited information cues regarding flagging procedures in all stages, it is unclear how users interpret them and what shapes their understanding. Therefore, we ask:
@@
RQ1: What are users’ mental models of how platforms implement flagging mechanisms, and how do they develop these models?
Next, we examine why users choose to flag content. Prior research on a gaming platform (Kou and Gui, 2021) shows that gamers often flag content for reasons that do not align with the platform’s notion of toxicity (e.g., flagging teammates’ poor game performance). We focus on flagging behaviors in social media platforms and the factors that drive users to flag by asking:
@@
RQ2: What motivates users to engage (or not engage) in flagging on social media sites?
Submitting a flag often involves traversing a multi-step interface that asks users to fit their objections with a post or account under the platforms’ pre-defined categories of inappropriate conduct. We explore the cognitive and privacy-related challenges of submitting a flag by asking:
@@
RQ3: What are users’ cognitive and security concerns about engaging in flagging?
@@
Applying interpretive qualitative analysis (Merriam, 2002) to our interview data, we found that the lack of transparency in flagging creates substantial gaps in users’ mental models about how platforms evaluate flagged content. However, users developed strategies to leverage their interpretation of how flags work, using flags alongside other site mechanisms, such as blocking and private messaging, to address online harms. A belief in generalized reciprocity motivates users to flag inappropriate content—they expect a collective flagging effort would help keep online spaces clean. However, flagging is not without its costs. We surfaced reservations users have about the cognitive burdens of flagging and privacy concerns when discussing potential innovations to flagging mechanisms.
@@
    Empirical data on user experiences and perspectives about flagging on social media platforms. We provide empirical data contributing to Crawford and Gillespie’s theoretical conceptualization of flagging  (Crawford and Gillespie, 2016). Our findings highlight the importance of flags on social media as a vital way for users to voice concerns within an opaque content moderation system. We explore diverse views of flags—as a right, obligation, exploitation, and a commons problem. We argue that platforms should promote digital literacy around flagging without imposing it as a duty.
@@
    Design opportunities with recommendations. We provide empirically informed design recommendations and identify opportunities for innovation. We recommend incorporating seams (Eslami et al., 2016) into flagging interfaces to help users understand their operations and ensure transparency throughout the flagging process. Innovations to flagging mechanisms should address user needs while respecting privacy and cognitive burden. Additionally, we identify opportunities for platforms to collaborate with marginalized communities to address the flagging-related challenges they face.
@@ 2.1. Content Moderation Affordances on Social Media
Most contemporary platforms hosting user-generated content offer a mechanism for users to provide feedback on the content they see that violates the platform’s guidelines, known as a “flag.” Platforms use flags to identify potentially inappropriate content, which human moderators or bots subsequently review  (Kou and Gui, 2021). If this review determines that the content violates the platform’s guidelines, the flagged user may face direct consequences, such as content removal and account suspension  (Jhaver et al., 2019a, 2021).
@@
We contextualize flags as part of a collection of mechanisms that let users express their voices in content moderation systems. These mechanisms commonly involve expressing dissatisfaction about types of content (e.g., hate speech, violence, spam)  (Kumar et al., 2021; Banko et al., 2020; Scheuerman et al., 2021) or behaviors (e.g., harassment, bullying, scamming)  (Shaw, 2018). Flags are often accompanied by a variety of personal moderation tools, which have been defined as “tools that let users configure or customize some aspects of their moderation preferences on social media”  (Jhaver et al., 2023). Personal moderation tools include account-based tools that let users mute or block specific accounts (Geiger, 2016; Matias et al., 2015) and content-based tools that let users hide comments containing configured keywords  (Jhaver et al., 2022) or set up content sensitivity controls  (Jhaver and Zhang, 2023). We focus on flagging, one of the most conspicuous content moderation mechanisms available to general end-users. While personal moderation tools change only the configuring user’s news feed, flags enable site-wide content removal, provide higher granularity in letting users express their dissatisfaction, and prompt users to justify their actions. Appeal mechanisms  (Vaccaro et al., 2020, 2021), which let moderated users report their dissatisfaction with content moderation decisions and request that they be reversed, are often used to contest platform moderation decisions about flagged content or other content removed by administrators. Users may also utilize the above moderation mechanisms in combination with regular platform interactions that demonstrate positive or negative algorithmic feedback (e.g., like and dislike buttons) (Khan, 2017) or other social signals (e.g., unfollowing accounts)  (Sasahara et al., 2021).
@@
Users with special roles (e.g., community moderators) usually have access to a wider range of moderation tools  (Hatmaker, 2021; Stockinger et al., 2023). For example, Reddit moderators may independently program Automod, a bot, to automatically identify and take action against content that infringes upon their community’s guidelines (Wright, 2022; Jhaver et al., 2019b). Similarly, Twitch streamers and YouTube creators rely on third-party applications for harmful content detection (Cai and Wohn, 2019; Jhaver et al., 2022). We do not focus on such specialized tools because this study is examining regular end-users who lack access to them. However, we draw from literature on accountability and division of labor involved in using such tools  (Jhaver et al., 2019b; Anaobi et al., 2023; Chandrasekharan et al., 2019) to examine how end users perceive the use of automation and human moderators to review flagged content.
@@ 2.2. User Perspectives on Content Moderation
Prior work asserted the importance of examining user perspectives of content moderation to inform platform leadership and policymakers. Scholars argue that these internet governance regimes are defining users’ roles and shaping our political and public discourse, so it is imperative to incorporate users’ perspectives in their design  (Riedl et al., 2021; Langvardt, 2017). Our research on users’ perceptions of flagging mechanisms highlights existing user needs to inform changes to how platforms design and administer these systems.
The discussion of user perspectives includes questions of who owns the responsibility for platform governance. Gillespie advocated for a collective civic responsibility to govern platforms  (Gillespie, 2018). However, it is not always clear how we may distinguish specific stakeholder responsibilities to apply this form of collective governance. For instance, users have indicated that they attribute responsibility for interventions and detecting inappropriate content to platforms (Riedl et al., 2021) or human moderators  (Naab et al., 2018) rather than themselves. However, users were found to prefer personally configurable moderation tools that regulate hate speech, violent content, and sexually explicit content instead of platform-enacted bans of those categories (Jhaver and Zhang, 2023). We build on this work by examining how end-users perceive their responsibilities to flag fitting in with how flagging responsibilities are divided.
@@
As an extension of discussions regarding the responsibilities to flag harmful content and the burden it places on users, it is known that people of marginalized identities are more likely to experience harm on online platforms (Duggan, 2017). For example, Black and female-identifying users were shown to disproportionately experience an array of harms from unwanted behaviors on platforms (e.g., hate speech, doxxing)  (Musgrave et al., 2022). To address these challenges, HCI and CSCW researchers have prioritized the content moderation needs and perspectives of vulnerable groups (Haimson et al., 2021; Thach et al., 2022, 2024; Liang et al., 2021). For example, Blackwell et al. examined HeartMob users’ experiences on a platform built for those most affected by severe online abuse  (Blackwell et al., 2017). They argue that platforms’ design and moderation must integrate vulnerable users’ unique needs to “fully” address online harassment. This is also the rationale for our emphasis on examining the diverse experiences of users from marginalized identities with flagging mechanisms. We explore participants’ concerns about flagging with a particular focus on fears of retaliation, concerns about privacy, and misuse of flagging mechanisms.
When considering the design of moderation systems, it is also important to consider end-users’ preferences for control and transparency. Currently, most social media platforms do not offer adequate visibility into the rules and the decision-making processes behind their moderation decisions  (Suzor et al., 2019). Sometimes, transparency challenges are a result of issues with existing moderation interfaces that, for example, do not allow for moderator transparency  (Juneja et al., 2020) or transparency about what decisions are made by automation or AI versus human moderators  (Ozanne et al., 2022). Some researchers have examined how different configurations of moderation tool design shape the control that users perceive over their social media feeds (Jhaver et al., 2022, 2023). Additionally, when moderation systems indicate the reasoning behind content regulation decisions, user attitude toward decisions may improve, resulting in higher-quality user posts  (Jhaver et al., 2019c, a; Kim, 2009) or generally increased trust  (Brunk et al., 2019). Flagging systems similarly lack transparency as platforms are often unclear about when they act on flagged content and what actions they perform. Such practices result in user perspectives being overlooked and unheard. For this reason, we offer empirical insights about how users view individual agency in flagging processes and examine their transparency needs.
@@ 2.3. Flagging Mechanisms on Social Media Platforms
While it is unclear when flags were first implemented on social media platforms, the notion of raising a concern to a community leader or administrator has long been used as a strategy for expressing dissatisfaction civilly. Flags are a crucial means for users to express such dissatisfaction in online environments.
@@
Previous research examined flagging from end-users’ perspectives as creators of content, especially in the context of how users develop “folk theories” (DeVito et al., 2017) about moderation decisions when platforms fail to provide adequate or any explanations for those decisions. The opaqueness of flagging has led to speculations about platforms making controversial decisions about flagged content without the flagged user’s knowledge, such as with shadowbanning (where platforms hide a user’s content from others’ feeds without removing it entirely)  (West, 2018). This has raised concerns about the use of flagging to silence (Are, 2023) or otherwise censor users (Peterson et al., 2013). Our study expands on this work by specifically examining the perspectives of users when they are flagging others’ content rather than being flagged themselves.
@@
The definition of how “flaggable” a post is, or its “flaggability,” for users may differ from platforms’ definitions of what is inappropriate. For example, users on gaming platforms utilized flags to report what they defined as toxic content but also flagged teammates to whom they attributed game losses (Kou and Gui, 2021). What’s more, what is interpreted as harmful enough to flag may differ from user to user (Jhaver et al., 2018a), and platforms have differing guidelines on what is deemed inappropriate and thus flaggable (Pater et al., 2016). These findings highlight a gap in our understanding of how users navigate conflicts between their values and platform guidelines for flagging and whether such appropriations of flags occur on social media platforms. Thus, we examine the decision-making process behind how users choose to flag inappropriate content online. Drawing from prior literature, we refer to inappropriate content as a multitude of behaviors (Scheuerman et al., 2021) that include but are not limited to trolling  (Gerlitz and Helmond, 2013; Phillips, 2015), invasion of privacy  (Sambasivan et al., 2019; Chen et al., 2023; Boyd, 2008; Aïmeur and Schőnfeld, 2011), public shaming  (Vaccaro et al., 2021; Ronson, 2015; Goldman, 2015), and interpersonal harm (Riedl et al., 2021).
At the moment, most major social media platforms provide users with the option to flag content. However, the actions users must take after hitting the initial button to flag content vary. For instance, when flagging content on YouTube, a user just clicks on the “report” button, selects a category of harm, and then is thanked for their efforts. At this point, users receive no notifications about the outcome of flagging such content. The flagged content may also continue to appear in their search results. Other platforms such as TikTok, X (previously Twitter), and Instagram offer the option for users to block or mute (i.e., not recommend the content of) accounts they flag or that are responsible for posting content they flagged. Facebook, at the moment, has the most objectively visible post-flagging process, providing users with a specific portal to check the status of their flags (Fig. 2). While previous studies have examined interventions similar to flags to mitigate online harm  (Katsaros et al., 2022; Chang and Danescu-Niculescu-Mizil, 2019; Jhaver et al., 2023), we expand this work in our focus on all three stages of the flagging process: 1) before 2) during and 3) after submitting a flag.

### v3

# Cleaning Up the Streets: Understanding Motivations, Mental Models, and Concerns of Users Flagging Social Media Content

@@ Abstract.
Social media platforms offer flagging, a technical feature that empowers users to report inappropriate posts or bad actors to reduce online harm. The deceptively simple flagging interfaces on nearly all major social media platforms disguise complex underlying interactions among users, algorithms, and moderators. Through interviewing 25 social media users with prior flagging experience, most of whom belong to marginalized groups, we examine end-users’ understanding of flagging procedures, explore the factors that motivate them to flag, and surface their cognitive and privacy concerns. We found that a lack of procedural transparency in flagging mechanisms creates gaps in users’ mental models, yet they strongly believe that platforms must provide flagging options. Our findings highlight how flags raise critical questions about distributing labor and responsibility between platforms and users for addressing online harm. We recommend innovations in the flagging design space that enhance user comprehension, ensure privacy, and reduce cognitive burdens.
@@ 1. Introduction
End-users have several options to take action when encountering inappropriate content on social media platforms, from regular interactions built into a platform (e.g., unfollowing an account) to content moderation-specific interactions. Of these options, it is only through flagging that end-users can directly request platform administrators to take site-wide actions against the content. Platforms maintain content regulation systems (comprising a coordinated deployment of automated tools and human reviewers (Jhaver et al., 2019b; Gorwa et al., 2020)) that regularly review flagged items and, when warranted, trigger sanctions, such as removing flagged posts or banning flagged accounts (Goldman, 2021; Grimmelmann, 2015). Thus, the flag serves as a powerful and empowering tool for user-driven moderation.
Though at first glance, flag implementations may signal platforms’ commitment to enacting democratic governance and fostering user communities, in practice, flags are often designed without adequately considering the needs of users who rely on them. As Crawford and Gillespie describe, flagging is a “complex interplay between users and platforms, humans and algorithms, and the social norms and regulatory structures of social media”  (Crawford and Gillespie, 2016). However, while their work identified the conceptual significance of flags as critical socio-technical mechanisms, there is little empirical research examining how users—the primary stakeholders in the flagging processes—experience and perceive these systems. This lack of focus on user perspectives is particularly important given that flagging processes across platforms significantly vary regarding the input required of flag submitters and post-flagging review and notification procedures.
@@
For instance, when a user flags a video on YouTube, they see a quick pop-up that thanks users for flagging content with no other contact from the platform about the outcome of their flag (Fig. 1). In contrast, when users flag content on Facebook, the platform typically tells them they will be notified when a decision is made. It provides them with additional safety options such as blocking, muting, and hiding the flagged content. Facebook also offers users a centralized interface where they may revisit past flags to check their review status (Fig. 2). These different implementations suggest that current flagging systems are frequently designed in an ad hoc fashion, and it remains unclear if they adequately understand user needs despite users being central to their success.
(a) The report button for flagging content on Youtube.
(b) Categories for flagging content on Youtube.
(c) Thank you pop-up after users flag content on Youtube.
Figure 1. YouTube’s flagging interface pages.
Figure 2. Facebook’s support inbox dashboard showing posts a user has flagged.
@@
In this study, we analyze user experiences with flagging mechanisms on major social media platforms and ask three key research questions. First, we note that flagging interfaces provide varying degrees of information regarding how the user input during flagging will be processed, how long it will take, who will review the flag, and more. It is unclear how users interpret these information cues and what shapes their understanding. Therefore, we ask:
@@
RQ1: What are users’ mental models of how social media platforms implement flagging mechanisms, and how do they develop these models?
Next, we examine why users choose to flag content on social media platforms. While flagging mechanisms have been studied in other contexts, such as gaming platforms, where users may flag for reasons unrelated to platform-defined toxicity (e.g., flagging teammates for poor game performance (Kou and Gui, 2021)), there is limited understanding of why users flag harmful content on social media. Social media platforms are unique in their role as central hubs for public discourse, making it critical to understand the factors driving users to flag or choose to not flag harmful content. Investigating these motivations is essential for ensuring that flagging mechanisms are designed to meet users’ needs and align with platform goals. Thus, we ask:
@@
RQ2: What motivates users to engage (or not engage) in flagging on social media platforms?
Submitting a flag is an act of anonymous and voluntary labor that requires users to navigate a multi-step interface, where they must categorize their concerns under the platform’s pre-defined labels of inappropriate conduct. This process not only demands cognitive effort to evaluate and classify content but also raises questions about users’ privacy and potential risks. To better understand these challenges, we investigate the cognitive and privacy-related implications of flagging by asking:
@@
RQ3: What are users’ cognitive and security concerns about engaging in flagging on social media platforms?
@@
Applying interpretive qualitative analysis (Merriam, 2002) to our interview data, we found that flagging a post involves three temporally distinct stages from the flagger’s perspective: (1) before, (2) during, and (3) after flagging. However, within each of these stages, the lack of transparency in flagging creates substantial gaps in users’ mental models about how platforms evaluate flagged content. Given this lack of transparency, users develop strategies to leverage their interpretation of how flags work, using flags alongside other site mechanisms, such as blocking and private messaging, to address online harms. A belief in generalized reciprocity motivates users to flag inappropriate content—they expect a collective flagging effort would help keep online spaces clean. However, flagging is not without its costs. We surface the reservations users have about the cognitive burdens of flagging and examine their privacy concerns when discussing potential innovations to flagging mechanisms.
@@
    Empirical data on user experiences and perspectives about flagging on social media platforms. We provide empirical data contributing to Crawford and Gillespie’s theoretical conceptualization of flagging  (Crawford and Gillespie, 2016). Our findings highlight the importance of flags on social media as a vital way for users to voice concerns within an opaque content moderation system. We also explore views of flags as a right and obligation, surfacing tensions between user motivations to flag and platform motives for offering flags. We provide insights about the labor of flagging.
@@
    Design opportunities with recommendations. We provide empirically informed design recommendations and identify opportunities for innovation in building flagging systems. We recommend incorporating seams (Eslami et al., 2016) into flagging interfaces to help users understand their operations and ensure transparency throughout the flagging process. We recommend that innovations to flagging mechanisms should address user needs while respecting privacy and cognitive burdens.
@@ 2.1. Content Moderation Affordances on Social Media
Most contemporary platforms hosting user-generated content offer a mechanism for users to provide feedback on the content they see that violates the platform’s guidelines. This mechanism is usually called “flag” or “report.” Platforms use flags to identify potentially inappropriate content, which human moderators or bots subsequently review  (Kou and Gui, 2021). If this review determines that the content indeed violates the platform’s guidelines, the flagged user may face direct consequences, such as content removal and account suspension  (Jhaver et al., 2019a, 2021).
@@
We contextualize flags as part of a collection of mechanisms that let users express their voice in content moderation systems. These mechanisms commonly involve expressing dissatisfaction about norm-violating content (e.g., hate speech, violence, spam)  (Kumar et al., 2021; Banko et al., 2020; Scheuerman et al., 2021) or behaviors (e.g., harassment, bullying, scamming)  (Shaw, 2018). Flags are often accompanied by a variety of personal moderation tools, which have been defined as “tools that let users configure or customize some aspects of their moderation preferences on social media”  (Jhaver et al., 2023). These personal moderation tools include account-based tools that let users mute or block specific accounts (Geiger, 2016; Matias et al., 2015; Jhaver et al., 2018b) and content-based tools that let users hide comments containing configured keywords  (Jhaver et al., 2022) or set up content sensitivity controls  (Jhaver and Zhang, 2023). We focus on flagging, one of the most conspicuous content moderation mechanisms available to general end-users. While personal moderation tools change only the configuring user’s news feed, flags enable site-wide content regulation, provide higher granularity in letting users express their dissatisfaction, and prompt users to justify their actions. Appeal mechanisms  (Vaccaro et al., 2020, 2021), which let moderated users report their dissatisfaction with content moderation decisions and request that they be reversed, are often used to contest platform moderation decisions about flagged content or other content removed by administrators. Users may also utilize the above moderation mechanisms in combination with regular platform interactions that demonstrate positive or negative feedback (e.g., like and dislike buttons) (Khan, 2017) or other social signals (e.g., unfollowing accounts)  (Sasahara et al., 2021).
@@
Users with special roles (e.g., community moderators) usually have access to a wider range of moderation tools  (Hatmaker, 2021; Stockinger et al., 2023). For example, Reddit moderators may independently program Automod, a bot, to automatically identify and take action against content that infringes upon their community’s guidelines (Wright, 2022; Jhaver et al., 2019b). Similarly, Twitch streamers and YouTube creators rely on third-party applications for harmful content detection (Cai and Wohn, 2019; Jhaver et al., 2022). We do not focus on such specialized tools because this study is examining regular end-users who lack access to them. However, we draw from literature on accountability and division of labor involved in the deployment of such tools  (Jhaver et al., 2019b; Anaobi et al., 2023; Chandrasekharan et al., 2019) to examine how end users perceive the use of automation and human moderators to review flagged content.
@@ 2.2. User Perspectives on Content Moderation
Prior work has asserted the importance of examining user perspectives of content moderation to inform platform practices and policymaking. Scholars argue that these internet governance regimes are defining users’ roles and shaping our political and public discourse, so it is imperative to incorporate users’ perspectives in their design  (Riedl et al., 2021; Langvardt, 2017). Our research on users’ perceptions of flagging mechanisms highlights existing user needs to inform changes to how platforms design and administer these systems.
The discussion of user perspectives includes questions about who owns the responsibility for platform governance. Gillespie advocated for a collective civic responsibility to govern platforms  (Gillespie, 2018). However, it is not always clear how we may distinguish specific stakeholder roles when applying this form of collective governance. For instance, prior research shows that users attribute the responsibility for interventions and detection of inappropriate content to platforms (Riedl et al., 2021) or human moderators  (Naab et al., 2018) rather than themselves. However,  Jhaver and Zhang (2023) showed that users prefer personally configurable moderation tools to regulate hate speech, violent content, and sexually explicit content instead of platform-enacted bans of those categories. This suggests that users desire to shape their feeds but expect platforms to intervene when necessary. We build on this work by examining how end-users perceive their responsibilities to flag fitting in with how they assume the obligations to regulate inappropriate content are divided among all stakeholders.
@@
As an extension of discussions regarding the responsibilities to flag harmful content and the burden it places on users, it is known that people of marginalized identities are more likely to experience harm on online platforms (Duggan, 2017). For example, Black and female-identifying users were shown to disproportionately experience an array of harms from unwanted behaviors on platforms (e.g., hate speech, doxxing)  (Musgrave et al., 2022). To address these challenges, HCI and CSCW researchers have prioritized the content moderation needs and perspectives of vulnerable groups (Haimson et al., 2021; Thach et al., 2024; Liang et al., 2021). For example, Blackwell et al. examined HeartMob users’ experiences on a platform built for those most affected by severe online abuse  (Blackwell et al., 2017). They argue that platforms’ design and moderation must integrate vulnerable users’ unique needs to “fully” address online harassment. This is also the rationale for our emphasis on examining the diverse experiences of users from marginalized identities with flagging mechanisms. We explore participants’ concerns about flagging with a particular focus on fears of retaliation, concerns about privacy, and misuse of flagging mechanisms.
When considering the design of moderation systems, it is also important to consider end-users’ preferences for control and transparency. Currently, most social media platforms do not offer adequate visibility into the rules and the decision-making processes behind their moderation decisions  (Suzor et al., 2019). Sometimes, transparency challenges are a result of issues with existing moderation interfaces that, for example, do not allow for transparency about what decisions are made by automated tools versus human moderators  (Ozanne et al., 2022; Juneja et al., 2020). Some researchers have examined how different configurations of moderation tool design shape the control that users perceive over their social media feeds (Jhaver et al., 2022, 2023). Additionally, prior research shows that when moderation systems indicate the reasoning behind content regulation decisions, user attitude toward those decisions may improve, resulting in higher-quality user posts  (Jhaver et al., 2019c, a; Kim, 2009) or generally increased trust  (Brunk et al., 2019). Flagging systems similarly lack transparency as platforms are often unclear about when they act on flagged content and what actions they perform. Such practices result in user perspectives being overlooked and unheard. For this reason, we examine users’ transparency needs and offer empirical insights about how they seek individual agency in flagging processes.
@@ 2.3. Flagging Mechanisms on Social Media Platforms
While it is unclear when flags were first implemented on social media platforms, the notion of raising a concern to a community leader or administrator has long been used to express dissatisfaction civilly (Dibbell, 1994). Flags are a crucial means for users to express such dissatisfaction in online environments. In this section, we provide an overview of related research on flagging as well as a description of flagging mechanisms on select social media platforms to better situate the context of our study.
@@
Previous research examined flagging from end-users’ perspectives as creators of sanctioned content, especially in the context of how users develop “folk theories” (DeVito et al., 2017) about moderation decisions when platforms fail to provide adequate or any explanations for those decisions (Thach et al., 2024). The opaqueness of flagging has led to speculations about platforms making controversial decisions about flagged content without the flagged user’s knowledge, such as with shadowbanning (where platforms hide a user’s content from others’ feeds without removing it entirely)  (West, 2018). This has raised concerns about the use of flagging to silence (Are, 2023) or otherwise censor users (Peterson et al., 2013). Our study expands on this work by specifically examining the perspectives of users when they are flagging others’ content rather than being flagged themselves.
@@
The definition of how “flaggable” a post is, or its “flaggability,” for users may differ from platforms’ definitions of what is inappropriate. For example, users on gaming platforms utilized flags to report what they defined as toxic content but also flagged teammates to whom they attributed game losses (Kou and Gui, 2021). Zhao and Chen (2023) reported a case study of mass reporting on Webo where fans of Chinese celebrities publicly coordinated to report content critical of those celebrities in an effort to take it down. What’s more, what is interpreted as harmful enough to flag may differ from user to user (Jhaver et al., 2018a; Luo and Li, 2024), and platforms have differing guidelines on what is deemed inappropriate and thus flaggable (Pater et al., 2016). These findings highlight a gap in our understanding of how users navigate conflicts between their values and platform guidelines for flagging. It also remains unclear how such appropriations of flags occur on social media platforms. Thus, we examine the decision-making process behind how users choose to flag inappropriate content online. Drawing from prior literature, we refer to inappropriate content as a multitude of behaviors (Scheuerman et al., 2021) that include but are not limited to trolling  (Gerlitz and Helmond, 2013; Phillips, 2015), invasion of privacy  (Sambasivan et al., 2019; Chen et al., 2023; Boyd, 2008; Aïmeur and Schőnfeld, 2011), public shaming  (Vaccaro et al., 2021; Ronson, 2015; Goldman, 2015), and interpersonal harm (Riedl et al., 2021).
Wang et al. (2023) examined users’ privacy concerns when reporting messages or accounts perpetrating harassment on end-to-end encrypted messaging platforms. They found that users make nuanced decisions about whether and how much to report based on their perceived trade-offs between privacy risks and protections. Further, these decisions are influenced by their trust in messaging platforms and community moderators. We build on this work to examine social media users’ privacy concerns when they consider reporting inappropriate content. Riedl and Newell (2024) reported a case study of Taylor Swift fans reporting nonconsensual pornographic deepfakes of the pop superstar on X. They noted that such reporting is motivated by fans’ strong sense of responsibility to protect Taylor Swift online. We add to this research by examining how users perceive their versus platforms’ flagging obligations more broadly.
At the moment, major social media platforms allow users to flag content, but post-flagging actions vary widely–from no updates on removal decisions (YouTube, Fig. 1) to automatically blocking accounts for users who flag them (X, formerly Twitter) and offering the option to unfollow or block accounts (Instagram and TikTok) to letting users track flagged content through dashboards (Facebook, Fig. 2). Despite the prevalence of flagging, research has primarily focused on other moderation strategies (Katsaros et al., 2022; Chang and Danescu-Niculescu-Mizil, 2019; Jhaver et al., 2023). Our study fills this gap by analyzing the user experience of flagging mechanisms across platforms, offering empirical insights to design flagging systems that better support users in addressing inappropriate content.